est the claim about the population mean μ at the level of significance α. Assume the population is normally distributed. Claim: μ > 28; α = 0.05; σ = 1.2 Sample statistics: = 28.3, n = 50
Answer #1

= z = Test rin solution Here Ho ll<28 vs Hiu>28 Given n=50 t=28-3, r=1.2 t-4 28.3-28 statistic 12/550 = 768 Given x=0.05, cri